20) If a Sample Size n is less than 5% of the Population Size N, then we must apply the Finite Population Correction Factor (FPCF) to the Standard Deviation of the Sample Proportion:  This Standard Deviation is then Given by

Find the Standard Deviation of the Sample Proportion if p = .64, N = 400, and n = 16.
Answer #1

Solution:

From given ,

n/N = 16/400 = 0.04 = 4% which is less than 5%

Population correction factor is applied when n is greater than 5% of the N.

So here , population correction factor is not applied.

The standard deviation of the sample proportions is

  =     

=

=  0.024